Calculate E tot for the electrochemical cells for A. Fe - Fe^+2 and Cu- CU

B. Zn - Zn^+2 and Ag- Ag
1
Expert's answer
2020-07-20T06:26:30-0400

A. E0(Fe/Fe2+) = -0,44 V = Еа

E0(Cu/Cu2+) = 0,34 V = Ес

E = Ec - Ea = 0,34 + 0,44 = 0,78 V

B. E0(Zn/Zn2+) = -0,763 V = Ea

E0(Ag/Ag+) = 0,799 V = Ec

E = Ec - Ea = 0,799 + 0,763 = 1,562 V

a = anode

c = cathode